OpenStudy (fanduekisses):

Am I doing this right? In the example we modeled the world population in the second half of the 20th century by the equation P(t) = 2560e^0.017185t. Use this equation to estimate the average world population during the time period of 1950 to 1980. (Round your answer to the nearest million.)

OpenStudy (fanduekisses):

So I have to integrate from 0 to 30 and divide by 30 right?

OpenStudy (fanduekisses):

\[\frac{ 1 }{ 30 } \int\limits_{0}^{30}2560e^{0.017185t} \]

OpenStudy (fanduekisses):

right

OpenStudy (peachpi):

yes

OpenStudy (fanduekisses):

I don't get the answer in millions though

OpenStudy (fanduekisses):

@ganeshie8

OpenStudy (will.h):

tell me if am wrong . that's an exponential function and the t= time however am not sure what e is!

OpenStudy (fanduekisses):

yes and e is just a popular number in math. I believe it's irrational. like pi

OpenStudy (will.h):

thats possible. so the function outlook is super confusing. is this the correct function \[p(t) 2560 \pi (0.017185)^t\] is that correct? if yes then all we need to do is to substitute the values of t and after that use the slope formula to determine the average.

OpenStudy (fanduekisses):

no, it's \[P(t)= 2560e^{0.017185t}\]

OpenStudy (will.h):

yeah so did you substitute the values 1950 and 1980 instead of t?

OpenStudy (fanduekisses):

but first, I have to integrate right?, because yes I did. well, 1950= 0 and 1980= 30

OpenStudy (will.h):

so there fore we have the coordinates (1950,0) and (1980,30) use slope formula y2-y1/x2-x1 to find the average.

OpenStudy (fanduekisses):

no, I meant that 1950 represents t=0 and 1980= 30 since 30 years past

OpenStudy (fanduekisses):

I'm pretty sure I have to use integration some how

zepdrix (zepdrix):

sure, integrate. are you confused because of the coefficient in the exponent?

OpenStudy (fanduekisses):

I'm going to show you guys how I do it.

OpenStudy (will.h):

please do

OpenStudy (fanduekisses):

\[\frac{ 1 }{ 30 }*2560\int\limits_{0}^{30}e^{0.017185t} dt\]

OpenStudy (fanduekisses):

\[\frac{ 1 }{ 30 }[2560(\frac{ e^{0.017185t} }{ 0.017185 })]_{0}^{30}\]

zepdrix (zepdrix):

Mmmm ok looks good! Then you just jam a bunch of numbers into your calculator, ya?

OpenStudy (fanduekisses):

lol but I get 8315.14

zepdrix (zepdrix):

Hmm I came up with 3349.6

OpenStudy (fanduekisses):

ohh wait a minute I forgo to subtract after plugging in the lower bound zero

zepdrix (zepdrix):

\[\large\rm 4965.57 e^{.017185}|_0^{30}\]If you bring all the coefficient junk to the front you have something like this. Ah yes, e^0 is not 0 :)

OpenStudy (fanduekisses):

yeah now I get 3349.57

zepdrix (zepdrix):

So the population averaged 3350 million people over that time period. Or 3.35 billion people.

OpenStudy (fanduekisses):

ohhh yay :)

OpenStudy (fanduekisses):

thanks!